### v1

# Daria Sushnikova    George Turkiyyah    Edmond Chow    David Keyes

Abstract
This paper presents a new fast iterative solver for large systems involving kernel matrices. Advantageous aspects of $\mathcal{H}^{2}$ matrix approximations and the multigrid method are hybridized to create the $\mathcal{H}^{2}$-MG algorithm. This combination provides the time and memory efficiency of $\mathcal{H}^{2}$ operator representation along with the rapid convergence of a multilevel method. We describe how $\mathcal{H}^{2}$-MG works, show its linear complexity, and demonstrate its effectiveness on two standard kernels. The current zoo of $\mathcal{H}^{2}$ solvers, which includes a wide variety of iterative and direct solvers, so far lacks a method that exploits multiple levels of resolution, commonly referred to in the iterative methods literature as “multigrid” from its origins in a hierarchy of grids used to discretize differential equations. This makes $\mathcal{H}^{2}$-MG a valuable addition to the collection of $\mathcal{H}^{2}$ solvers. The algorithm has potential for advancing various fields that require the solution of large, dense, symmetric positive definite matrices.
keywords:
$\mathcal{H}^{2}$matrix, multigrid methods, kernel matrices, rank-structured matrices, iterative solvers, linear complexity
{MSCcodes}
65F10, 65N55, 65F30, 65F55
1 Introduction
This paper tackles the challenge of solving linear systems with large, dense kernel matrices. Such systems arise in a wide range of applications, including computational statistics [3, 38], machine learning [31, 29], and computational physics [4, 10]. Solving these systems is particularly challenging due to their quadratic and cubic complexity in terms of memory and runtime, respectively. Over the past decades, significant progress has been made to address this issue through rank-structured matrix approximations.
Rank-structured methods, and particularly $\mathcal{H}^{2}$  [20, 5] matrices, typically provide a time- and memory-efficient matrix-by-vector product, which naturally leads to solving systems using iterative solvers. However, iterative solvers have their disadvantages, as their efficiency depends on the number of iterations and thus on the matrix conditioning. Direct solvers have their own challenges, being extremely complex for rank-structured matrices and involving a large constant overhead. To fill this gap, we looked at the multigrid [12, 6, 17] method, which is typically used for sparse matrices, and adapted it for $\mathcal{H}^{2}$ matrices. Multigrid methods exhibit excellent convergence properties and can significantly benefit from fast $\mathcal{H}^{2}$ matrix-vector products. In this paper, we introduce a new algorithm, $\mathcal{H}^{2}$- multigrid ($\mathcal{H}^{2}$-MG ), which leverages the hierarchical structure of $\mathcal{H}^{2}$ matrices to create a multigrid method that operates across different levels of the $\mathcal{H}^{2}$ matrix.
The key contributions of this work include:
  • Developing the $\mathcal{H}^{2}$-MG algorithm, a hybrid solver combining multilevel resolution with hierarchical matrix representation.
  • Theoretically demonstrating the linear complexity of the algorithm in both time and memory.
  • Validating its effectiveness on two standard kernels and comparing its performance with existing approaches.
The proposed method merges the time and memory efficiency of $\mathcal{H}^{2}$ matrices with the fast convergence properties of multigrid. This paper not only expands the repertoire of $\mathcal{H}^{2}$ solvers but also addresses a critical gap in the literature by providing a multigrid-inspired approach to hierarchical matrix methods. The simplicity and scalability of the $\mathcal{H}^{2}$-MG algorithm make it a valuable addition to the field, with potential applications across diverse domains.
2 Related work
A general $N$ by $N$ matrix requires $\mathcal{O}(N^{2})$ operations to compute a matrix-vector product and $\mathcal{O}(N^{2})$ memory for storage. However, significant progress has been made to reduce this computational cost with controllable loss of accuracy. Block low-rank methods, such as the mosaic skeleton [35, 36] and the $\mathcal{H}$-matrix [21, 18, 25] techniques, reduce the number of operations for matrix-vector products and storage to $\mathcal{O}(N\log N)$ by exploiting low-rank approximations of certain blocks of the matrix. Building upon these advancements, nested-basis methods like HSS matrices [39, 8, 13], HODLR matrices [1], $\mathcal{H}^{2}$ matrices [20, 5, 27], and the Fast Multipole Method (FMM) [15, 40, 14] were proposed. These methods enable matrix-vector multiplication with high accuracy in $\mathcal{O}(N)$ operations for many matrices arising in physically causal models where interactions decay smoothly with distance. HSS and HODLR methods are particularly efficient for 1D problems, while $\mathcal{H}^{2}$ and FMM extend their efficiency to 2D and 3D problems thanks to their strong admissibility property.
This accelerated matrix-vector multiplication becomes the basis for solving linear systems using iterative techniques such as GMRES [30], CG [22, 33], BiCGstab [37], and any other iterative solvers. Iterative methods, while versatile, have a drawback: their convergence rate depends on the conditioning or eigenvalue clustering of the matrix. In contrast, fast direct solvers guarantee a predefined number of operations for solving the system. Although the complexity for general dense matrices remains $\mathcal{O}(N^{3})$, leveraging the $\mathcal{H}$-matrix format has led to breakthroughs in direct solver efficiency. Researchers in [19] introduced an $\mathcal{O}(N\log N)$ direct solver algorithm, albeit with a substantial constant. Subsequent works [2, 28, 34] utilizing $\mathcal{H}^{2}$ (FMM) format have achieved $\mathcal{O}(N)$ direct solver algorithms with more favorable constants.
In this paper, we expand the zoo of $\mathcal{H}^{2}$ solvers by introducing a novel iterative algorithm – the $\mathcal{H}^{2}$-MG solver. This solver is rooted in the standard multigrid method but tailored for $\mathcal{H}^{2}$ structures.
Multigrid (MG) methods, as well as rank structured methods, play a crucial role in solving large linear systems, particularly systems with sparse matrices arising from discretized partial differential equations. Multigrid has its origins in [12] and was further developed in [6]. Multigrid has become a cornerstone of numerical techniques for solving large sparse linear systems. Its strategy lies in efficiently reducing errors at multiple scales by leveraging a hierarchy of coarser grids. Over time, a variety of fruitful generalizations have emerged, such as algebraic multigrid (AMG) [7] approaches that extend its applicability to unstructured grids and irregular geometries [11, 9, 16] and fast multipole preconditioners for sparse matrices [23].
Our approach of applying multigrid to the system with $\mathcal{H}^{2}$ matrix diversifies the landscape of $\mathcal{H}^{2}$ solvers, offering a promising alternative that is relatively easier to implement and parallelize efficiently compared to direct solvers. Demonstration of parallel scaling, however, is beyond the scope of this initial description.
3 Algorithm
3.1 $\mathcal{H}^{2}$ matrix
In this section, we briefly review the fundamental concept of an $\mathcal{H}^{2}$ matrix, a hierarchical block low-rank matrix structure with a nested basis property. Matrices well approximated in $\mathcal{H}^{2}$ form typically come from the discretization of boundary integral equations and several other problems with approximately separable kernels. The hierarchical nature of $\mathcal{H}^{2}$ matrix is the basis and main inspiration of the $\mathcal{H}^{2}$-MG algorithm. For more comprehensive and formal $\mathcal{H}^{2}$ definitions, see [15, 20, 5].
Consider the linear system
$Ax=b,$
where $A\in\mathbb{R}^{N\times N}$, is dense and $x,b\in\mathbb{R}^{N}$. Let the rows and columns of matrix $A$ be partitioned into $M$ blocks. The size of $i$-th block is $B_{i}$, $i\in 1,\dots.,M$. Each block $A_{ij}$, $i,j\in 1,\dots.,M$, of matrix $A$ has either full rank, denoted by
$A_{ij}=\widehat{D}_{ij},\quad\widehat{D}_{ij}\in\mathbb{R}^{B_{i},B_{j}},$
and is called a “close” block, or has a low rank, is called “far”, and possess the following property:
(1) $A_{ij}\approx\widehat{F}_{ij}=\widehat{U}_{i}\widehat{S}_{ij}\widehat{V}_{j},$
where $\widehat{D}_{ij}\in\mathbb{R}^{B_{i}\times B_{j}}$, $\widehat{U}_{i}\in\mathbb{R}^{B_{i}\times r_{i}}$, $\widehat{S}_{ij}\in\mathbb{R}^{r_{i}\times r_{j}}$, $\widehat{V}_{j}\in\mathbb{R}^{r_{j}\times B_{j}}$. Note that all low-rank blocks in a row $i$ have the same left factor $\widehat{U}_{i}$, and all the low-rank blocks in a column $j$ have the same right factor $\widehat{V}_{j}$. This is one of the defining features of the $\mathcal{H}^{2}$ matrix. We denote $r_{i}$ as the rank of $i$-th block row, excluding full-rank blocks, and $r_{j}$ is the rank of $j$-th block column, excluding full-rank blocks.
Let us define a block matrix $D\in\mathbb{R}^{N\times N}$:
$[D]_{ij}=\begin{cases}\widehat{D}_{ij},&\quad\text{if }A_{ij}\text{ is a closeblock}\\ 0,&\quad\text{if }A_{ij}\text{ is a far block}\end{cases}.$
Note that $D$ is typically a block-sparse matrix; it contains a number of nonzero blocks per block row that is independent of dimension. Also, define a block matrix $F\in\mathbb{R}^{N\times N}$:
$[F]_{ij}=\begin{cases}0,&\quad\text{if }A_{ij}\text{ is a close block}\\ \widehat{F}_{ij}=\widehat{U}_{i}\widehat{S}_{ij}\widehat{V}_{j},&\quad\text{if}A_{ij}\text{ is a far block}\end{cases}.$
The matrix $A$ is split into two matrices:
(2) $A=D+F,$
To write equation (1) in a matrix form, we define rectangular diagonal matrices $U_{1}\in\mathbb{R}^{N\times N_{2}}$, $V_{1}\in\mathbb{R}^{N_{2}\times N}$, where ${N_{2}=\sum_{i=1}^{M}r_{i}}$.
$U_{1}=\begin{bmatrix}\widehat{U}_{1}&&\\ &\ddots&\\ &&\widehat{U}_{M}\end{bmatrix},\quad V_{1}=\begin{bmatrix}\widehat{V}_{11}&&\\ &\ddots&\\ &&\widehat{V}_{MM}\end{bmatrix},$
Also, define a matrix $S_{1}\in\mathbb{R}^{N_{2}\times N_{2}}$:
$[S]_{ij}=\begin{cases}0,&\quad\text{if }A_{ij}\text{ is a close block}\\ \widehat{S}_{ij},&\quad\text{if }A_{ij}\text{ is a far block}\end{cases}.$
These definitions allow us to rewrite equation (1) in a matrix form:
$F=U_{1}S_{1}V_{1}$
The equation (2) then evolves to:
$A=D+U_{1}S_{1}V_{1}.$
This decomposition is illustrated in Figure 1.
[…]
Figure 1: Illustration of the one-level $\mathcal{H}^{2}$ matrix
The figure illustrates a special case; the matrix $D$ may have more complex block structure. In general, it could be any block-sparse matrix.
The matrix $S_{1}$ is a dense matrix. Our next goal is to compress it using the same strategy as we applied to matrix $A$. We assemble together blocks of matrix $S_{1}$ in patches of size $p$. (Typical choices of $p$ are $2$ or $2^{d}$, where $d$ is the physical dimension of the problem.) We obtain matrix $S_{1}$ with $\frac{M}{p}$ blocks. Those blocks again separate into a set of full-rank close and low-rank far blocks. Analogously to the previous procedure, we obtain the factorization
(3) $S_{1}=D_{2}+U_{2}S_{2}V_{2},$
where $D_{2}\in\mathbb{R}^{N_{2}\times N_{2}}$, $U_{2}\in\mathbb{R}^{N_{2}\times N_{3}}$, $V_{2}\in\mathbb{R}^{N_{3}\times N_{2}}$, $S_{2}\in\mathbb{R}^{N_{3}\times N_{3}}$. The size of $N_{3}$ is $\sum_{i=1}^{\frac{M}{p}}r_{i}$, where $r_{i}$ are the ranks of the far blocks of the matrix $S_{1}$. An illustration of the factorization (3) is in Figure 2.
[…]
Figure 2: Block low-rank factorization of the matrix $S_{1}$
The process continues until the matrix $S_{l}$ on the $l^{th}$ level has low-rank blocks. Assuming $l=2$, we obtain:
$A=D+U_{1}\left(D_{2}+U_{2}S_{2}V_{2}\right)V_{1}.$
In the general case:
(4) $A=D+U_{1}\left(D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2}\right)V_{1}.$
This recursive summation is called the $\mathcal{H}^{2}$ approximation of the matrix $A$.
Multiplication of a vector $x\in\mathbb{R}^{N}$ by the $\mathcal{H}^{2}$ matrix $A\in\mathbb{R}^{N\times N}$ follows the formula (4):
$y=Ax=Dx+U_{1}\left(D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2}\right)V_{1}x.$
Denote $x_{i}=V_{i}x_{i-1}$, $x_{i}\in\mathbb{R}^{N_{i}}$, $i=1\dots l$, assuming $x=x_{0}$, to obtain:
$y=Dx+U_{1}\left(S_{1}x_{1}+U_{2}(\dots(S_{l-1}x_{l}+U_{l}S_{l}x_{l})\dots)\right).$
Denote $y_{i-1}=D_{i}x_{i-1}+U_{i}y_{i}$, $i=1,\dots,l$, $y_{l}=S_{l}x_{l}$, and note that that $D_{1}=D$, $y_{0}=y$, to obtain:
$y=Dx+U_{1}y_{1}.$
Figure 3 shows a schematic of the procedure.
[…]
Figure 3: Schematic of $\mathcal{H}^{2}$ matrix-by vector multiplication, $l=2$
Remark 1.
For the $\mathcal{H}^{2}$ matrix $A\in\mathbb{R}^{N\times N}$, according to [5], the storage requirement scales as $\mathcal{O}(N)$. $\mathcal{H}^{2}$ construction and matrix-by-vector multiplication complexity is also $\mathcal{O}(N)$.
3.2 Multigrid method
A multigrid method [6, 17] is a powerful numerical technique used for solving partial differential equations (PDEs) and linear systems of equations with sparse matrices. While a solver in its own right, it is often used to precondition other iterative algorithms to cause them to converge rapidly.
The essential idea of multigrid is to define and solve problems on multiple grids of varying levels of resolution, referred to as the multigrid hierarchy. These grids range from coarse to fine, with each level representing a discretization of the problem at a different scale. At the coarsest level, the problem is smaller and easier to solve. Solutions obtained at this level are then recursively interpolated to finer grids. This hierarchical approach allows the method to address errors more effectively, each on an optimal scale, and accelerate convergence.
The multigrid method operates in cycles, typically consisting of phases of smoothing and inter grid transfer (restriction and prolongation). In “smoothing”, iterative methods like Jacobi [24], Gauss-Seidel [32], CG [22, 33], etc., are applied to reduce errors. (Smoothing is in quotation marks, because this operation can sometimes “roughen” the solution.) In restriction, information is passed from finer to coarser grids and, in prolongation, from coarser to finer grids.
In this paper, as is often done in other contexts, we broaden the idea of multigrid beyond problems arising from discretized PDEs, here to systems with $\mathcal{H}^{2}$ matrices. Analogously to using operators on multiple grid levels in a standard approach, we use the levels of $\mathcal{H}^{2}$ hierarchy, shown in Figure 3. The transfer matrices $U_{i}$ and $V_{i}$, $i\in 1\dots l$ in $\mathcal{H}^{2}$ serve naturally as restriction and prolongation operators. The derivation of a multilevel iteration applied to the $\mathcal{H}^{2}$ system ($\mathcal{H}^{2}$-MG) is detailed in Section 3.3.
One of the key advantages of multigrid is its ability to rapidly converge to a highly accurate solution, often achieving convergence rates that are essentially independent of the problem size. This makes its translation to systems with $\mathcal{H}^{2}$ matrices highly beneficial and leads to a new and effective solver for systems with $\mathcal{H}^{2}$ matrices.
3.3 $\mathcal{H}^{2}$-MG
In this section, we implement a multigrid algorithm for a system with $\mathcal{H}^{2}$ matrix. Consider the system
(5) $Ax=b,$
where $A\in\mathbb{R}^{N\times N}$ is an $l$-level $\mathcal{H}^{2}$ matrix, $b\in\mathbb{R}^{N}$ is the right-hand side, $x_{0}\in\mathbb{R}^{N}$ is the initial guess for the solution.
Let us follow the steps of the classic multigrid method, applying them to the system (5). First, we reformulate the system in terms of error and residual. Let the residual of the system be
$r_{1}=b-Ax_{0}.$
Then, subtracting from both parts of equation (5) the $Ax_{0}$ term we write:
$Ax-Ax_{0}=b-Ax_{0}.$
Noting that $Ax-Ax_{0}=A(x-x_{0})=Ae_{1}$, where $e_{1}=x-x_{0}$ is the error, we obtain the error residual equation:
(6) $Ae_{1}=r_{1},$
as an equivalent form of  (5).
The next stage of multigrid is the application of smoothing iterations to the system (5). We consider several iterations of some iterative method on the system (6) as a smoother, $\textbf{Iter}(A,r_{1},0)$. In this notation, the first parameter is the operator, the second is the right-hand side, the third is the initial guess, and the result is the solution after the iterations.
In our numerical experiments, which are on symmetric positive definite matrices, we use CG [33, 22] solver as a smoother because of its fast and easily parallelized application. Thanks to the $\mathcal{H}^{2}$ structure of matrix $A$, the application of one iteration of CG is linear in time and memory cost. After smoothing, we receive an approximation of the error:
$\tilde{e}_{1}=\textbf{Iter}(A,r_{1},0).$
Then, we compute the residual:
$\hat{r}_{1}=r_{1}-A\tilde{e}_{1}.$
Then we subtract the term $A\tilde{e}_{1}$ from both parts of equation (6),
$Ae_{1}-A\tilde{e}_{1}=r_{1}-A\tilde{e}_{1}.$
Denoting $\hat{e}_{1}=e_{1}-\tilde{e}_{1}$ to obtain the system
(7) $A\hat{e}_{1}=\hat{r}_{1}.$
Now is the time to build the restriction and prolongation operators. If we denote the matrix $A$ explicitly in its $\mathcal{H}^{2}$ format, we obtain an expanded version of the equation (7):
$D+U_{1}\left(D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2}\right)V_{1}\hat{e}_{1}=\hat{r}_{1},$
from which we see that a straightforward way to restrict the system is to multiply it by $U^{\top}_{1}$ on the left and to insert matrix $I=V^{\top}_{1}V_{1}$ between $A$ and $x$. We obtain the restricted system:
$U^{\top}_{1}(D+U_{1}\left(D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2}\right)V_{1})V^{\top}_{1}V_{1}\hat{e}_{1}=U^{\top}_{1}\hat{r}_{1},$
or, if we open the first parentheses:
$(U^{\top}_{1}DV^{\top}_{1}+U^{\top}_{1}U_{1}\left(D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2}\right)V_{1}V^{\top}_{1})V_{1}\hat{e}_{1}=U^{\top}_{1}\hat{r}_{1}.$
Then, reduce all identity matrices:
$(U^{\top}_{1}DV^{\top}_{1}+D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2})V_{1}\hat{e}_{1}=U^{\top}_{1}\hat{r}_{1}.$
We denote the restricted operator by $A_{2}$:
(8) $A_{2}=U^{\top}_{1}DV^{\top}_{1}+D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2}).$
Remark 2.
Note that $A_{2}\in\mathbb{R}^{N_{2}\times N_{2}}$ has $\mathcal{H}^{2}$ structure, just like matrix $A$. The matrices $A$ and $A_{2}$ have exactly the same structure, except that matrix $A$ stores the full matrix $D$, while $A_{2}$ stores a reduced part of it, $U_{1}^{\top}DV^{\top}$. Thus, matrix $A_{2}$ stores less information than $A$. Since $A$ scales linearly as an $\mathcal{H}^{2}$ matrix, $A_{2}$ also scales linearly. See the detailed proof in Section 3.4.
Denote the restricted error and residual vectors:
$e_{2}=V_{1}\hat{e}_{1},$
$r_{2}=U^{\top}_{1}\hat{r}_{1}.$
In our case, the $U^{\top}_{i}$ matrices are the analogs of the restriction operators, $V^{\top}_{i}$ are the analogs of the prolongation operators, and the basis vectors of the $\mathcal{H}^{2}$ levels are the analogs of the coarser grids. We obtain the restricted system:
$A_{2}e_{2}=r_{2}$
Then, according to the multigrid algorithm, we apply smoother to the restricted system and obtain the approximation of the error:
$\tilde{e}_{2}=\textbf{Iter}(A_{2},r_{2},0).$
This continues until we reach the top level $l$. At this level, we have a system
$A_{l}e_{l}=r_{l},$
where $A_{l}$ is a small dense matrix since $l$ is the top level of $\mathcal{H}^{2}$ hierarchy. We solve the system directly with dir_sol. In our computations, we use the Cholesky factorization:
$e_{l}=\textbf{dir\_sol}(A_{l},r_{l}).$
Then, we move back from coarser to finer grids. We apply the prolongation operator $V_{l}^{\top}$ to the error $e_{l}$ and correct the $\tilde{e}_{l-1}$ error:
$\tilde{e}_{l-1}=\tilde{e}_{l-1}+V_{l}^{\top}e_{l}.$
Then, we apply the smoothing operator starting with initial guess $\tilde{e}_{l-1}$:
$e_{l-1}=\textbf{Iter}(A_{l-1},r_{l-1},\tilde{e}{l-1}).$
We continue until we reach level 1. From the estimated error $e_{1}$, we obtain the approximation of the solution $x^{*}$:
$x^{*}=x_{0}+e_{1}.$
This is analogous to the multigrid V-cycle. We can perform multiple V-cycles to obtain a more accurate solution, using $x^{*}$ as the initial guess for the next V-cycle. Figure 4 is a visualization of $\mathcal{H}^{2}$-MG V-cycle.
$\tilde{e}_{1}=\textbf{Iter}(A,r1,0)$$e_{1}=\textbf{Iter}(A,r_{1},\tilde{e}_{1})$$r_{2}=U_{1}^{\top}(r_{1}-A\tilde{e}_{1})$$r_{3}=U_{2}^{\top}(r_{2}-A_{2}\tilde{e}_{2})$$\tilde{e}_{1}=\tilde{e}_{1}+V_{1}^{\top}e_{2}$$\tilde{e}_{2}=\tilde{e}_{2}+V_{2}^{\top}e_{3}$$\tilde{e}_{2}=\textbf{Iter}(A_{2},r_{2},0)$$e_{2}=\textbf{Iter}(A_{2},r_{2},\tilde{e}_{2})$$e_{3}=\textbf{dir\_sol}(A_{3},r_{3})$
Figure 4: Schematic of the $\mathcal{H}^{2}$-MG algorithm
Remark 3.
The visualization of $\mathcal{H}^{2}$-MG V-cycle, presented in Figure 4, is flipped upside-down to emphasize the analogy of $\mathcal{H}^{2}$-MG method with the $\mathcal{H}^{2}$ structure. Compare $\mathcal{H}^{2}$-MG V-cycle in Figure 4 and the $\mathcal{H}^{2}$ matrix by vector product in Figure 3.
In Algorithm 1, we give the formal description of one V-cycle of the $\mathcal{H}^{2}$-MG algorithm. $A=A_{1}$, matrices $A_{i}\in\mathbb{R}^{N_{i}\times N_{i}}$, $i=1,\dots,l-1$, are the $\mathcal{H}^{2}$ matrices, matrix $A_{l}\in\mathbb{R}^{N_{l}\times N_{l}}$ is dense, $x_{0}\in\mathbb{R}^{N}$ is the initial guess, and $b\in\mathbb{R}^{N}$ is a right-hand side.
Algorithm 1 One V-cycle of the $\mathcal{H}^{2}$-MG algorithm
1:$r_{1}=b-A_{1}x_{0}$
[…]
3:     $\tilde{e}_{i}=\textbf{Iter}(A_{i},r_{i},0)$
4:     $r_{i+1}=U^{\top}_{i}(r_{i}-A_{i}\tilde{e}_{i})$
5:end for
6:$e_{l}=\textbf{dir\_sol}(A_{l},r_{l})$
[…]
8:     $\tilde{e}_{i}=\tilde{e}_{i}+V^{\top}_{i}e_{i+1}$
9:     $e_{i}=\textbf{Iter}(A_{i},r_{i},\tilde{e}_{i})$
10:end for
11:$x^{*}=x_{0}+e_{1}$
We can run successive V-cycles using the output $x^{*}$ as the next initial guess. It may be of interest to consider multigrid cycles beyond V-cycles, namely W- or F-cycles.
Remark 4.
In $\mathcal{H}^{2}$-MG , we treat the number of smoothing iterations on the finest level (with matrix $A$) and on the coarser levels (with matrices $A_{i}$, $i=2,\dots,l$) as two separate parameters. We denote the number of fine-level iterations as $n_{f}$ and the number of coarse-level iterations as $n_{c}$. Unlike the standard multigrid method, in our case, we have a physical grid only at the finest level, while all other grids are “basis” grids of the $\mathcal{H}^{2}$  matrix. Therefore, we treat the finest level of the problem differently by assigning it an independent number of smoothing iterations. In the numerical section, we empirically demonstrate that this approach leads to better convergence.
3.4 Complexity analysis
In this subsection, we describe the time and memory complexity of the $\mathcal{H}^{2}$-MG algorithm.
The crucial feature of the $\mathcal{H}^{2}$-MG complexity analysis is the linear complexity of the $\mathcal{H}^{2}$ matrix. According to Remark 1, for the $\mathcal{H}^{2}$ matrix of size $N\times N$, the construction complexity, the memory requirements, and matrix-by-vector product complexity are all $\mathcal{O}(N)$. Assume $c_{H2}$ to be the $\mathcal{H}^{2}$ matrix-by-vector product constant.
First, compute the $\mathcal{H}^{2}$-MG computational complexity $n_{\text{op}}$ to run one V-cycle. The complexity of the fine grid smoothing iterations is $n_{f}c_{H2}N$, the coarse grid smoothing iterations is $n_{c}c_{H2}N_{i}$, for $i=2,\dots,l-1$, the complexity of the direct solver is $c_{d}N_{l}^{3}$, where $c_{d}$ is the direct solver complexity constant, and the restriction and prolongation operator complexity is negligible, compared to the smoothing iterations. The overall complexity is:
$n_{\text{op}}=2n_{f}c_{H2}N+2n_{c}\sum_{i=2}^{l-1}c_{H2}N_{i}+c_{d}N_{l}^{3}.$
Assume, for simplicity, that the block size is $B$ for all blocks and block rank $r$ is fixed for all levels (consider it to be the maximum rank of any block). Also, assume the number of blocks of the initial matrix $A$ is $M$, and the number of blocks stacked together while transferring from level to level is $d$. Thus, $N=MB$, $N_{i}=\frac{Mr}{d^{i-2}}$, for $i=2,\dots,l$, and the overall complexity is:
$n_{\text{op}}=2n_{f}c_{H2}MB+2n_{c}\sum_{i=2}^{l-1}\frac{Mrc_{H2}}{d^{i-2}}+c_{d}N_{l}^{3}.$
Using the sum of geometric progression, we obtain:
$n_{\text{op}}=2n_{f}c_{H2}MB+2n_{c}\left(\frac{d-\frac{1}{d^{l-3}}}{d-1}\right)c_{H2}Mr+c_{d}N_{l}^{3}.$
By the construction of $\mathcal{H}^{2}$ , the size of the coarsest level $N_{l}$ is a constant; thus,
$n_{\text{op}}=\mathcal{O}(N),$
with the constant
$c_{\text{op}}=2n_{f}c_{H2}+2n_{c}\left(\frac{d-\frac{1}{d^{l-3}}}{d-1}\right)c_{H2}\frac{r}{B}$
Now compute the memory requirements. The $\mathcal{H}^{2}$-MG algorithm requires storage of matrices $A$, $A_{i}$ for $i=2,\dots,l$, $U_{i}$, and $V_{i}$ for $i=1,\dots,l-1$. Matrices $U_{i}$ and $V_{i}$ are block-diagonal, and their storage is negligible compared to matrices $A$ and $A_{i}$. Assume $\mathcal{H}^{2}$ memory constant to be $c_{m}$. Thus, the memory requirements are:
$n_{\text{mem}}=c_{m}N+\sum_{i=2}^{l-1}c_{m}N_{i}+N_{l}^{2}.$
Analogously to the time complexity rewriting matrix sizes in terms of $M$, $B$, $r$, and $d$ and applying the sum of geometric progression, we obtain:
$n_{\text{mem}}=c_{m}N+\left(\frac{d-\frac{1}{d^{l-3}}}{d-1}\right)c_{m}Mr+N_{l}^{2}.$
Taking into account that $N_{l}$ is a constant, obtain:
$n_{\text{mem}}=\mathcal{O}(N),$
with the overall $\mathcal{H}^{2}$-MG storage constant
$c_{\text{mem}}=c_{m}+\left(\frac{d-\frac{1}{d^{l-3}}}{d-1}\right)c_{m}\frac{r}{B}.$
Thus, one V-cycle of the $\mathcal{H}^{2}$-MG algorithm is linear in both time and memory, though its constant factor is larger compared to a single iteration of the CG algorithm. However, in practice, we observe that the number of $\mathcal{H}^{2}$-MG V-cycles needed to achieve the required accuracy is independent of problem size, whereas the number of CG iterations needed to reach the same accuracy increases with the problem size.
4 Numerical results
The numerical experiments showcase our Python implementation of the $\mathcal{H}^{2}$-MG algorithm, demonstrating $\mathcal{H}^{2}$-MG convergence for two different kernels. The matrices are cast into the $\mathcal{H}^{2}$ format using the MCBH [27, 26] method. The code for the $\mathcal{H}^{2}$-MG algorithm is publicly available at https://github.com/dsushnikova/h2mg.
We compare the $\mathcal{H}^{2}$-MG algorithm with solvers CG and FMM-LU [34]. All three methods benefit from $\mathcal{H}^{2}$ structure of matrix $A$; thus, the comparison is natural. The $\mathcal{H}^{2}$-MG algorithm uses CG iteration for smoothing, illustrating how the coarse-level iterations improve the convergence relative to CG alone. In the plots, we present convergence of the $A$-norm of the error $||e_{k}||_{A}$ on $k$-th iteration. We consider convergence per iteration and per time. Note that the plotted timings include the computation of the $A$-norms. For the CG method we plot all the iterations, for the $\mathcal{H}^{2}$-MG we plot only the fine grid iterations (with the matrix $A$), the coarser levels iterations are seen on the convergence plots as error drops.
In all our tests, we randomly generate $x_{\text{true}}$, then compute $b=Ax_{\text{true}}$ and solve the system $Ax=b$ for $x$.
Remark 5.
The right-hand sides were chosen in this manner so that the error norm at each iteration could be easily computed. We also tested a few cases where the right-hand sides were chosen randomly from a standard Gaussian distribution. In these cases, the linear systems were harder to solve, but H2-MG outperformed unpreconditioned CG, just like in the detailed results that we show below.
4.1 Gaussian kernel
For a first example, we consider the system
(9) $Ax=b,$
with the Gaussian matrix $A$. We consider a uniform tensor grid on a unit square $P\subset\mathbb{R}^{2}$: $p_{i}\in P$, $i\in 1\dots N$, where $N$ is the number of points. The kernel matrix $A$ is given by the formula:
(10) $a_{ij}=\begin{cases}\exp{(-\frac{|p_{i}-p_{j}|^{2}}{\sigma})},&\quad\text{if }i\neq j\\ 1+c,&\quad\text{if }i=j\end{cases},$
where $\sigma\in\mathbb{R}$ is the dispersion parameter, $c\in\mathbb{R}$ is a small regularization parameter. Matrix $A$ is approximated to the $\mathcal{H}^{2}$ format with accuracy $\epsilon=10^{-9}$ ($\epsilon$ is a relative error between the result of a matrix-vector multiplication performed using the $\mathcal{H}^{2}$ matrix and that using the original matrix), and the number of levels is chosen adaptively.
4.1.1 Gaussian matrix, analysis of error evolution
In our tests, we randomly generate $x_{\text{true}}$, then compute $b=Ax_{\text{true}}$ and solve the system $Ax=b$ for $x$. During the process, we can compute the error $e_{1*}=x_{k}-x_{\text{true}}$ since we know $x_{\text{true}}$. In the classic multigrid algorithm, the error $e_{1*}$ should behave in a following way: for the fine-grid iterations, the high-frequency components of $e_{1*}$ should decrease rapidly, and during the coarse-grid iterations, successive lower frequency components should decrease in turn.
In this section, we study the behavior of the error $e_{1*}$ in the $\mathcal{H}^{2}$-MG method. Consider $U_{1}$, the transfer matrix from the finest level to the coarser one. $U_{1}$ has orthonormal columns. Assume that the matrix $Q_{1}$ has columns that span the complimentary space.
To analyze the components of the error after smoothing in the basis of the interpolation matrix $U_{1}$ or in other words the part of the error that will be projected to the coarser levels, we compute $\text{pinv}(U_{1})e_{1*}$. Similarly, for the part of the after-smoothing error in the basis of the matrix $Q_{1}$, $\text{pinv}(Q_{1})e_{1e}$.
We study 1D Gaussian matrix with points equally spaced between 00 and $1$, $N=1024$, $\sigma=0.01$, $c=10^{-3}$. The results of this analysis are presented in Figure 5. The blue curves represent the error components after the first smoothing step in the initial $V$-cycle, the red curves are for the second V-cycle.
(a) Component of the post-smoothing error in the basis of $U_{1}$
(b) Component of the post-smoothing error in the basis of $Q_{1}$
Figure 5: Convergence comparison for different numbers of coarse iterations for the $\mathcal{H}^{2}$-MG method
The graphs clearly demonstrate that the error component $\text{pinv}(U_{1})e_{1*}$ is larger than that in the error component $\text{pinv}(Q_{1})e_{1*}$ components, providing compelling evidence that the CG smoothing can complement coarse grid correction effectively.
4.1.2 Gaussian matrix, analysis of number of smoothing steps
Consider the system (9) with matrix $A$ given by (10) with $c=10^{-3}$, $\sigma=0.1$. In Remark 4, we considered assigning the finest level of the problem a different number of smoothing steps than the coarser levels. In this subsection, we analyze the effect of parameters $n_{f}$ and $n_{c}$ on the $\mathcal{H}^{2}$-MG convergence.
Let us first study the convergence rate of the $\mathcal{H}^{2}$-MG depending on the number of fine iterations. Figure 6 presents the convergence comparison for different numbers of fine iterations, $n_{f}$ to the accuracy $\varepsilon=10^{-9}$, where $\varepsilon=\frac{(eAe^{\top})^{\frac{1}{2}}}{||b||_{2}}<10^{-9}$. The number of coarse iterations is fixed $n_{c}=40$. Fine and coarse iteration comparisons are given for matrix size $N=8\times 10^{4}$.
(a) $A$-norm per iteration
(b) $A$-norm per time
Figure 6: Convergence comparison for different numbers of fine iterations for the $\mathcal{H}^{2}$-MG method
We can see that in terms of convergence per iteration, fewer fine-grid iterations lead to better convergence. Figure 7 presents the convergence comparison to the accuracy $\varepsilon=10^{-9}$ for different numbers of coarse iterations, $n_{c}$. The number of fine iterations is fixed at $n_{f}=1$.
(a) $A$-norm per iteration
(b) $A$-norm per time
Figure 7: Convergence comparison for different numbers of coarse iterations for the $\mathcal{H}^{2}$-MG method
In this case, the number of coarse iterations directly influences the speed of convergence: the more coarse iterations, the faster the convergence. However, in terms of time, the optimal number of coarse iterations is an intermediate value ($n_{c}=40$). This is because there is a trade-off between the time spent on additional coarse iterations and the resulting improvement in convergence speed.
Remark 6.
Unlike a traditional multigrid method, where all levels are similar and require the same number of smoothing iterations, $\mathcal{H}^{2}$-MG uses a different number of smoothing iterations for the fine and coarse grids. The finest level is unique, as it is based on the physical grid, while all other levels are basis-induced, requiring a different number of smoothing steps.
4.1.3 Gaussian matrix, asymptotics analysis
In this subsection we consider the asymptotic behavior of the system (9) with matrix $A$ given by (10) for various combinations of parameters $c$ and $\sigma$. Figures 8 and 9 show the convergence of the $\mathcal{H}^{2}$-MG algorithm compared to CG across various problem sizes. The method tolerance is $\varepsilon=10^{-9}$. For the test we use parameters $n_{f}=1$, $n_{c}=40$.
(a) $c=10^{-3}$
(b) $c=10^{-5}$
Figure 8: Convergence evolution of $\mathcal{H}^{2}$-MG and CG as problem size increases, $\sigma=0.1$
(a) $c=10^{-3}$
(b) $c=10^{-5}$
Figure 9: Convergence evolution of $\mathcal{H}^{2}$-MG and CG as problem size increases, $\sigma=0.01$
We observe that strong regularization parameter $c=10^{-3}$ leads to the fast convergence of the $\mathcal{H}^{2}$-MG method for both matrices with $\sigma=0.1$ and $\sigma=0.01$, while the weak regularization parameter $c=10^{-5}$ leads to the divergence of both algorithms for the larger problem sizes.
The numbers of V-cycles required for convergence to a fixed accuracy are presented in Table 1.
Matrix Parameters Problem Size
1e4 2e4 4e4 8e4 16e4 32e4
$\sigma=0.1$, $c=10^{-3}$ 2 2 2 2 2 3
$\sigma=0.1$, $c=10^{-5}$ 4 4 3 5 4 7
$\sigma=0.01$, $c=10^{-3}$ 7 5 7 8 11 12
$\sigma=0.01$, $c=10^{-5}$ 200 85 156 235 428 711
Table 1: Number of V-cycles to solve the system for different problem sizes
As in the previous example, the number of V-cycles required for the convergence to the fixed accuracy does not grow significantly.
Figures 8 and 11 show the comparison of the overall solution time for the $\mathcal{H}^{2}$-MG and CG methods, compared against the H2-direct solver FMM-LU [34]. The goal of the comparison is to explore the efficiency of the proposed method by benchmarking it against an alternative efficient $\mathcal{H}^{2}$ solver. A red cross indicates that either the iterative method failed to converge within 5000 iterations, or the direct solver failed to solve the system with the required accuracy.
(a) $c=10^{-3}$
(b) $c=10^{-5}$
Figure 10: Overall solution time of methods $\mathcal{H}^{2}$-MG and CG, $\sigma=0.1$
(a) $c=10^{-3}$
(b) $c=10^{-5}$
Figure 11: Overall solution time of methods $\mathcal{H}^{2}$-MG and CG, $\sigma=0.01$
For $\sigma=0.01$ and $c=10^{-5}$, the system appears to have an extremely large condition number, leading to superlinear scaling or failure to converge within 5000 iterations of the iterative solvers. For the direct solver, this condition results in failure due to computational errors, as LU methods without pivoting struggle to handle systems with extremely ill-conditioned matrices. For three other cases, both $\mathcal{H}^{2}$-MG and FMM-LU solve the system with linear scaling. Since both $\mathcal{H}^{2}$-MG and FMM-LU scale linearly in this example, the main competition lies in the constant factor, where the iterative $\mathcal{H}^{2}$-MG naturally outperforms direct solver FMM-LU. $\mathcal{H}^{2}$-MG also outperforms CG in constant.
4.2 Exponential kernel
Consider the linear system (9) where $b\in\mathbb{R}^{N}$ is a right-hand side vector, $x\in\mathbb{R}^{N}$ is an unknown vector, and $A\in\mathbb{R}^{N\times N}$ is a kernel matrix with the linear exponential decay kernel. Consider a uniform tensor grid on a unit square $P\subset\mathbb{R}^{2}$: $p_{i}\in P$, $i\in 1\dots N$, where $N$ is the number of points. The kernel matrix $A$ is given by the formula:
(11) $a_{ij}=\begin{cases}\exp({-\frac{|p_{i}-p_{j}|}{\sigma}}),&\quad\text{if }i\neq j\\ 1+c,&\quad\text{if }i=j\end{cases},$
where $\sigma\in\mathbb{R}$ is the dispersion parameter of the matrix, $c\in\mathbb{R}$ is a constant. Matrix $A$ is approximated to the $\mathcal{H}^{2}$ format with accuracy $\epsilon=10^{-9}$, number of levels is chosen adaptively.
We keep all the assumptions made for the problem in Section 4.1 for this example.
4.2.1 Exponent matrix, asymptotics analysis
In this subsection we consider the asymptotic behavior of the system (9) with matrix $A$ given by (11) for various combinations of parameters $c$ and $\sigma$. Figures 12 and 13 show the convergence of the $\mathcal{H}^{2}$-MG algorithm compared to CG across various problem sizes. The method tolerance is $\varepsilon=10^{-9}$. We use parameters $n_{f}=1$, $n_{c}=40$.
(a) $c=10^{-3}$
(b) $c=10^{-5}$
Figure 12: Convergence evolution of $\mathcal{H}^{2}$-MG and CG as problem size increases, $\sigma=0.1$
(a) $c=10^{-3}$
(b) $c=10^{-5}$
Figure 13: Convergence evolution of $\mathcal{H}^{2}$-MG and CG as problem size increases, $\sigma=0.01$
We can see that strong regularization parameter $c=10^{-3}$ leads to the fast convergence of the $\mathcal{H}^{2}$-MG method for both matrices with $\sigma=0.1$ and $\sigma=0.01$, while the weak regularization parameter $c=10^{-5}$ leads to the divergence of both algorithms for the larger problem sizes.
The numbers of V-cycles required for the convergence to the fixed accuracy are presented in Table 2.
Matrix Parameters Problem Size
1e4 2e4 4e4 8e4 16e4 32e4
$\sigma=0.1$, $c=10^{-3}$ 14 13 18 25 54 128
$\sigma=0.1$, $c=10^{-5}$ 31 44 127 300 - -
$\sigma=0.01$, $c=10^{-3}$ 19 13 25 8 19 10
$\sigma=0.01$, $c=10^{-5}$ 22 15 30 14 30 18
Table 2: Number of V-cycles to solve the system for different problem sizes
As in the previous example, the number of V-cycles required for the convergence to the fixed accuracy does not grow significantly.
Figures 12(a) and 15 show the comparison of the overall solution time for the $\mathcal{H}^{2}$-MG and CG methods, compared against the H2-direct solver FMM-LU. A red cross indicates that either the iterative method failed to converge within 5000 iterations, or the direct solver failed to solve the system with the required accuracy.
(a) $c=10^{-3}$
(b) $c=10^{-5}$
Figure 14: Overall solution time of methods $\mathcal{H}^{2}$-MG and CG, $\sigma=0.1$
(a) $c=10^{-3}$
(b) $c=10^{-5}$
Figure 15: Overall solution time of methods $\mathcal{H}^{2}$-MG and CG, $\sigma=0.01$
For $\sigma=0.1$, the system appears to have an extremely large condition number, leading to quadratic scaling or failure to converge within 5000 iterations of the iterative solvers. For the direct solver, this condition results in failure due to computational errors, as LU methods without pivoting struggle to handle systems with extremely ill-conditioned matrices. For $\sigma=0.01$, both $\mathcal{H}^{2}$-MG and FMM-LU solve the system with linear scaling, while CG exhibits a higher scaling power of $y=x^{1.5}$. Since both $\mathcal{H}^{2}$-MG and FMM-LU scale linearly in this example, the main competition lies in the constant factor, where the iterative $\mathcal{H}^{2}$-MG naturally outperforms.
5 Conclusion
The $\mathcal{H}^{2}$-MG algorithm offers an advance in solving large, dense kernel matrices efficiently by iterative means. By combining the rapid convergence of the multigrid method with the time and memory efficiencies of $\mathcal{H}^{2}$ matrix approximations, this algorithm not only fills a gap in the existing suite of $\mathcal{H}^{2}$ solvers but also expands the toolkit available for tackling complex computational problems. The demonstrated linear complexity and practical effectiveness of $\mathcal{H}^{2}$-MG , verified through numerical examples, underscore its potential as a for applications burdened by large, dense, kernel matrices. Future work will aim to expand its applicability beyond symmetric positive definite matrices, high-performance implementation, and integration into other computational frameworks.